OpenStudy (anonymous):

Jim has 6.75 in dimes and quarters he had 8 more dimes than quarters how many coins of each type

hero (hero):

.25x + .10y = 6.75 y = x + 8

hero (hero):

y = (6.75 - .25x)/.10 y = x + 8 y = y .10(x+8) = 6.75 - .25x .10x + .8 = 6.75 - .25x .35x = 5.95 x = 5.95/.35 x = 17 y = 25 17(.25) + 25(.10) = 6.75

hero (hero):

Good luck figuring that out

OpenStudy (anonymous):

what do you mean?

hero (hero):

I'm sure you will have no problem understanding my steps.

OpenStudy (anonymous):

I am working on it.

hero (hero):

Okay

OpenStudy (anonymous):

If there is something helpful you can clue me in on please do

hero (hero):

It's pretty straight-forward from my point of view. Is there a step you're stuck on?

OpenStudy (anonymous):

I will have to continue to look at it. thanks

hero (hero):

Hint: at one point while solving, you cross-multiply. Good luck.

OpenStudy (anonymous):

sounds more like a riddle than trying to help me understand. I will have my tutor help me on this one, but thanks for your effort.